HUBBLE SPACE TELESCOPE DAILY REPORT #2610 PERIOD COVERED: 0000Z (UTC) 04/25/00 - 0000Z (UTC) 04/26/00 Daily Status Report as of 117/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 8439 (Bias Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no reported problems. 1.2 Completed Three Sets of WF/PC-2 8460 (Cycle 8 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations completed with no reported problems. 1.3 Two Sets of Completed STIS/CCD 8437 (Dark Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. The proposal completed nominally. 1.4 Completed FGS 8527 (Near Term Stability) FGS-2R was used to determine the near term stability of the instrument. The goals are twofold: {1} to establish the baseline Long Term Stability measurement for the new FGS and monitor changes in the relative plate scale and distortion since the SMOV MiniOFAD; and {2} to monitor changes in the transfer function over months. The proposal completed nominally. 1.5 Completed STIS/CCD 8803 (Focus Calibration Program) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the HST focus. The HST focus drifts slowly and perhaps undergoes slips of a few microns at random times. This focus monitoring program can provide evidence of a likely problem with the focus, but does not provide sufficient data upon which to act in a short period of time. This two orbit program obtains a large amount of high signal to noise focus data sufficient for accurately defining the HST focus. The program completed nominally. 1.6 Completed STIS/CCD/MA1 8241 (A Snapshot Survey of Interstellar Absorption Lines) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make high resolution Echelle spectrographic observations of the standard ultraviolet star HD212791 as part of a snapshot survey of interstellar absorption lines. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Three Sets of WF/PC-2 8805 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the WFPC2 Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working. The proposal completed with no reported problems. 1.8 Completed STIS/MA1 8426 (Cycle 8 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a routine Cycle-8 MAMA dark noise calibration measurement. This proposal will provide the primary means of checking on the health of the MAMA detectors. This is done through frequent monitoring of the background count rate. The observations were completed as planned, and no problems were reported. 1.9 Completed STIS/CCD 8789 (Probing the Outskirts of an Extrasolar Planet: Photometry and Spectroscopy) The Space telescope Imaging Spectrograph (CCD) was used to observe HD 209458, the first extrasolar planet known to transit the disk of its parent star, providing unique probes of this system for moons, rings and atmospheric composition that are unlikely to be feasible with ground-based observations. The proposal completed normally, with no reported problems. 1.10 Completed WF/PC-2 8240 (Masses and Spins of Black Holes in Seyfert Galaxies) The WF/PC-2 was used to observe about 20 bright Seyfert 1 galaxies that are now known to have very broad {up to 100, 000 km s^-1} Fe KAlpha emission lines near 6.4 keV. The observations completed nominally. 1.11 Completed Three Sets of STIS/CCD 8549 (POMS Test Proposal: Non-Scripted Parallel Proposal Continuation I) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. The observations were completed as planned, and no anomalies were reported. 1.12 Completed WF/PC-2 8234 (Calibrating Star Formation with the Metal-Rich Starburst Galaxy M83 {NGC5236}) The WF/PC-2 was used to study the structure of the prime nearby example of a metal-rich, actively star- forming environment: the starburst located in the nuclear region of the spiral galaxy M83. In M83 the PC gives 1 pc resolution so we can chart the detailed structure of this dusty but UV-bright system. For this purpose we will obtain broad-band images in the mid-UV, V, and I bands to locate and characterize stars, star clusters, and obscuring dust clouds, as well as HAlpha, HBeta, OIII, and SII narrow band images to diagnose conditions in the ionized ISM. The observations completed nominally. 1.13 Completed Two Sets of WF/PC-2 8445 (WF/PC-2 Cycle 8 Earth Flats) The WF/PC-2 was used to make Cycle 8 Earth flat calibration observations. This is to obtain sequences of earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed STIS/CCD 8796 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation IV) The Space Telescope Imaging Spectrograph (CCD) was used in the parallel mode to make some low galactic latitude archive observations. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed WF/PC-2 8146 (A Search for Binary L-Dwarfs) The WF/PC-2 was used to search for close binaries amongst the sample of 37 low-temperature L dwarfs, identified primarily in the course of the DENIS and 2MASS near-infrared surveys. The observations completed with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 6 Successful: 6 Scheduled Re-acquisitions: 9 Successful: 9 2.2 FHST Updates: Scheduled: 15 Successful: 15 2.3 Operations Notes: Using ROP IC-1, the SSA-2R transmitter was turned on and off for three intervals. Per ROP IC-1, MA transmitter #1 was turned on and off once. The SSR EDAC error counter was cleared at 116/1320Z per ROP SR-1A. The Engineering Status Buffer was dumped and cleared at 116/1330Z as directed by ROP DF-18A. Real-time operations transferred to CCS string "B" at 116/1441Z for another test of CCS Release 3.2 commanding. After the successful conclusion of the test, operations returned to the "A" string at 116/1456Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW